OpenStudy (dj3strella):

help me please 2. Where is the vertex of the parabola? y = x2 + 2x – 3 Above the x-axis Below the x-axis On the x-axis On the y-axis

OpenStudy (anonymous):

no habla engles

OpenStudy (anonymous):

I didn't learn that yet, ;( sorry

OpenStudy (anonymous):

que nina's

OpenStudy (dj3strella):

su ok

OpenStudy (dj3strella):

esta bien @SexyDestiny1

OpenStudy (anonymous):

ok lo siento

OpenStudy (anonymous):

el nina mensaje que me @DJ3strella

OpenStudy (anonymous):

y = x2 + 2x – 3 if we 'complete the square' we can pull out the coordinates for the vertex. step 1, take the coefficeint of the x term out, 2 step 2, half it 2/2 = 1 step 3, square it 1^2 = 1 step 4, add 1 and subtract one to the equation y = x2 + 2x – 3 y = x2 + 2x +1 – 3 -1 step 5, group the terms y = x2 + 2x +1 – 3 -1 y = (x2 + 2x +1) – 3 -1 step 6, simpify y= (x+1)^2 -4 step 7, extract coordiantes of vetex (h, k) y= (x+1)^2 -4 y = a(x-h)^2 - k can you see which values are 'h' and 'k'?

OpenStudy (dj3strella):

no @DemolisionWolf

OpenStudy (anonymous):

thats ok! there are a lot of steps in this, what parts don't make sense in 'completing the square'?